Lonnie is considering the purchase of a rental property with several units. The property rents for $8,600 a month when all units are occupied. When all units are occupied, additional income from on-sight laundry facilities is expected to be $200 a month. The units are expected to be rented 85% of the year. Additional expenses associated with the property include real estate taxes of $10,000 a year, liability insurance of $3,500 a year, advertising expense of $1,500 a year, maintenance costs of $12,500 a year, depreciation of $20,500 a year, and interest expense on the property loan of $24,000 a year. If Lonnie's required rate of return on the property is 11%, what is the intrinsic value of the property?

Select one:

a. $62,260.

b. $347,800.

c. $547,455.

d. $566,000.
Answer #1

Intrinsic value of property = Net operating income / Capitalisation rate

Net operating income = Earning from property - Opertaing expenses related to property

Operating expenses doent not inclue depreciation

Earning from Property = (8600+200)*12*85% [ 85% is done since unit occupied for 85% of year]

=89760

Operating expenses = Property tax +Insurance+Advertising expenses+ Maintenance cost +Interest expenses

= 10000+3500+1500+12500+24000 i.e. 51500

Net operating income =89760-51500 i.e. 38260

Assuming Net operating income for perpetuity

Intrinsic value = 38260/0.11 i.e. 347818 approx i.e. Option B